# Characterisation of a quantum bus between two driven qubits

### v2

@@ IV.2 Fidelity of the two-qubit gate
In the previous section we study how temperature affects the qubit-qubit entanglement generation, which is crucial for universal quantum computation. The thermal occupation of the system is determined by the temperature of its environment. As a rule of thumb, higher temperatures are detrimental to coherence. The system parameters, such as characteristic frequencies and coupling constants, are also critical in achieving high-fidelity quantum computing. Typically, these parameters are fixed by external electric or magnetic fields, materials choices, or the physical realisation of the quantum register. In the following, we study how different parameters of the model affect the fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate.
Fidelity $\mathcal{F}$ is a measure of how closely a quantum gate’s implementation matches the ideal behaviour ($\mathcal{F}=1$). For pure states, it is given by the inner product of the real and ideal final states, averaged over all initial pure states. However, thermal fluctuations and damping drive the system into a mixed state, so the definition of the gate fidelity takes a more general form:
@@
Fidelity is non-monotonous in $\Delta_{R}$. On the one hand, the rotating wave approximation necessary for building the harmonic oscillator-mediated $\sqrt{i\mathrm{SWAP}}$-gate [see Supplementary note A [29]], requires a small Rabi detuning $|\Delta_{R}|\ll\omega_{h}$. On the other hand, the dispersive regime is valid for $g\ll\Delta_{R}$. Focusing on $\Delta=0$ (blue line), the rotating wave approximation is well satisfied for Rabi detunings as large as $|\Delta_{R}|\approx 0.1\omega_{h}$, but the fidelity has a significant drop for $|\Delta_{R}|/g<10$ due to deviations from the dispersive regime. Regarding the driving field detuning, a nonvanishing $\Delta$ changes the qubits’ eigenbasis. Therefore, a high-fidelity gate implementation requires that $\Delta\ll\Omega_{R},\omega_{h}$. Based on the results of Fig. 4(a), an appropriate range of values for the detunings is $\Delta<0.01\omega_{h}$ and $10g<|\Delta_{R}|\lesssim 0.1\omega_{h}$. For a given $\Delta$, $|\Delta_{R}|$ may be calibrated to maximise $\mathcal{F}$.
@@
as depicted in Fig. 4(c). In Fig. 4(d) we show how the gate fidelity depends on the rise time. Based on the previous analysis of the system parameters, we select values of the detunings and coupling constant suitable for quantum computation. Modern quantum computers achieve two-qubit gate fidelities of approximately $\mathcal{F}\sim 99\%$ [44, 45, 46]. A minimum fidelity of $\mathcal{F}\sim 99.99\%$ [47, 48], combined with quantum error correction protocols, is expected for practical quantum computing. The general trend shows that gate fidelity decreases with slower rise times, but there is also an oscillatory behaviour that may be exploited to enhance $\mathcal{F}$. Given the practical limitations on the rise time, one may calibrate the rise time to maximise the fidelity.
V Practical realisation
Most present-day quantum computer realisations rely on qubits coupled via electromagnetic resonators or with other qubits connected to each other electromagnetically [32]. Such interqubit coupling requires that the qubit states are discernible based on their different charge susceptibilities. Those qubits are inherently susceptible to offset charge fluctuations, even if the effective coupling to those fluctuations can be minimised as done, for example, in superconducting transmon qubits [49]. The longest single-qubit coherence times are therefore obtained with qubits having no coupling to charge, such as spin qubits [50]. Then, the problem is to find a suitable mediator for coupling such qubits while maintaining their good coherence. One possible choice is based on utilizing mechanical vibrations in a mechanical resonator, coupling to the qubit states via any mechanism where the resonator’s motion shifts the (Zeeman) spin splitting, such as the qubit strain-induced effects [51], or a motion-modulated external magnetic field. The latter mechanism hence requires a spatial magnetic field gradient, and the coupling strength is proportional to the amount of change in the magnetic field within the zero-point vibration amplitude.
@@ VI Conclusions
In this work, we thoroughly analyze the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. In many experimental realisations, the characteristic frequencies of the bare qubits and the bus differ by several orders of magnitude. In such cases, dressing the qubits via a driving field provides a mechanism for selectively coupling the qubits to the bus. First, we investigate a qubit readout mechanism based on the measurement of the resonance frequency of the harmonic oscillator [33, 34, 16]. We show that the qubit state readout mechanism is more robust for the dressed qubit than the bare qubit as the steady state in the former generally exhibits information on both initial qubit states, and it is sufficient to fix the measurement time by observing the behaviour of the different amplitudes of the peaks. In contrast, for bare qubits, the resonance peak describing the steady state may mask the peak associated with the initial transient state. Moreover, dressed qubits tolerate thermal occupations better. We have also studied how the qubit-harmonic oscillator coupling can be leveraged to implement a two-qubit gate. We have analyzed the influence of various model parameters on the gate fidelity, identifying an optimal parameter regime for quantum computation.
@@
One key feature of using dressed qubits is the balanced steady state. A bare qubit is strongly affected by environmental damping, leading to relaxation into its ground state $|0\rangle$. In contrast, a dressed qubit relaxes into a mixed state, symmetrizing the $|0\rangle$ and $|1\rangle$ logical states. The state readout time may exceed the qubit’s decoherence time, so having a mixed steady state helps prevent a qubit flip due to thermal relaxation, thus improving the efficiency of initial state determination. However, most quantum algorithms require pure states, making the steady state of a dressed qubit unsuitable for qubit initialisation. A potential solution is to suppress the dressing amplitude $\Omega_{R}$ and apply a Zeeman field along the $z$-direction, enabling the qubit to decay into its bare ground state via thermal relaxation.
@@ Acknowledgements.
Methods
Code availability
The underlying code for this study is available in GitLab and can be accessed via this link https://gitlab.jyu.fi/jyucmt/quantum-bus-between-driven-qubits.
Data availability
The datasets generated and analysed during the current study are available in the GitLab repository https://gitlab.jyu.fi/jyucmt/quantum-bus-between-driven-qubits.
Author contributions
AH developed the code, generated the results, and conducted the analysis presented in this work. All authors contributed to the discussion of the results and the writing of the manuscript.
Competing interests
All authors declare no financial or non-financial competing interests.

### v3

@@ IV.2 Fidelity of the two-qubit gate
In the previous section we study how temperature affects the qubit-qubit entanglement generation, which is crucial for universal quantum computation. The thermal occupation of the system is determined by the temperature of its environment. As a rule of thumb, higher temperatures are detrimental to coherence. The system parameters, such as characteristic frequencies and coupling constants, are also critical in achieving high-fidelity quantum computing. Typically, these parameters are fixed by external electric or magnetic fields, materials choices, or the physical realization of the quantum register. In the following, we study how different parameters of the model affect the fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate.
Fidelity $\mathcal{F}$ is a measure of how closely a quantum gate’s implementation matches the ideal behavior ($\mathcal{F}=1$). For pure states, it is given by the inner product of the real and ideal final states, averaged over all initial pure states. However, thermal fluctuations and damping drive the system into a mixed state, so the definition of the gate fidelity takes a more general form:
@@
Fidelity is non-monotonous in $\Delta_{R}$. On the one hand, the rotating wave approximation necessary for building the harmonic oscillator-mediated $\sqrt{i\mathrm{SWAP}}$-gate [see Appendix A], requires a small Rabi detuning $|\Delta_{R}|\ll\omega_{h}$. On the other hand, the dispersive regime is valid for $g\ll\Delta_{R}$. Focusing on $\Delta=0$ (blue line), the rotating wave approximation is well satisfied for Rabi detunings as large as $|\Delta_{R}|\approx 0.1\omega_{h}$, but the fidelity has a significant drop for $|\Delta_{R}|/g<10$ due to deviations from the dispersive regime. Regarding the driving field detuning, a nonvanishing $\Delta$ changes the qubits’ eigenbasis. Therefore, a high-fidelity gate implementation requires that $\Delta\ll\Omega_{R},\omega_{h}$. Based on the results of Fig. 4(a), an appropriate range of values for the detunings is $\Delta<0.01\omega_{h}$ and $10g<|\Delta_{R}|\lesssim 0.1\omega_{h}$. For a given $\Delta$, $|\Delta_{R}|$ may be calibrated to maximize $\mathcal{F}$.
@@
as depicted in Fig. 4(c). In Fig. 4(d) we show how the gate fidelity depends on the rise time. Based on the previous analysis of the system parameters, we select values of the detunings and coupling constant suitable for quantum computation. Modern quantum computers achieve two-qubit gate fidelities of approximately $\mathcal{F}\sim 99\%$ [43, 44, 45]. A minimum fidelity of $\mathcal{F}\sim 99.99\%$ [46, 47], combined with quantum error correction protocols, is expected for practical quantum computing. The general trend shows that gate fidelity decreases with slower rise times, but there is also an oscillatory behavior that may be exploited to enhance $\mathcal{F}$. Given the practical limitations on the rise time, one may calibrate the rise time to maximize the fidelity.
V Practical realization
Most present-day quantum computer realizations rely on qubits coupled via electromagnetic resonators or with other qubits connected to each other electromagnetically [31]. Such interqubit coupling requires that the qubit states are discernible based on their different charge susceptibilities. Those qubits are inherently susceptible to offset charge fluctuations, even if the effective coupling to those fluctuations can be minimized as done, for example, in superconducting transmon qubits [48]. The longest single-qubit coherence times are therefore obtained with qubits having no coupling to charge, such as spin qubits [49]. Then, the problem is to find a suitable mediator for coupling such qubits while maintaining their good coherence. One possible choice is based on utilizing mechanical vibrations in a mechanical resonator, coupling to the qubit states via any mechanism where the resonator’s motion shifts the (Zeeman) spin splitting, such as the qubit strain-induced effects [50], or a motion-modulated external magnetic field. The latter mechanism hence requires a spatial magnetic field gradient, and the coupling strength is proportional to the amount of change in the magnetic field within the zero-point vibration amplitude.
@@ VI Conclusions
In this work, we thoroughly analyze the use of driven qubits coupled to a harmonic oscillator to implement a $\sqrt{i\mathrm{SWAP}}$-gate. In many experimental realizations, the characteristic frequencies of the bare qubits and the bus differ by several orders of magnitude. In such cases, dressing the qubits via a driving field provides a mechanism for selectively coupling the qubits to the bus. First, we investigate a qubit readout mechanism based on the measurement of the resonance frequency of the harmonic oscillator [32, 33, 16]. We show that the qubit state readout mechanism is more robust for the dressed qubit than the bare qubit as the steady state in the former generally exhibits information on both initial qubit states, and it is sufficient to fix the measurement time by observing the behavior of the different amplitudes of the peaks. In contrast, for bare qubits, the resonance peak describing the steady state may mask the peak associated with the initial transient state. Moreover, dressed qubits tolerate thermal occupations better. We have also studied how the qubit-harmonic oscillator coupling can be leveraged to implement a two-qubit gate. We have analyzed the influence of various model parameters on the gate fidelity, identifying an optimal parameter regime for quantum computation.
@@
One key feature of using dressed qubits is the balanced steady state. A bare qubit is strongly affected by environmental damping, leading to relaxation into its ground state $|0\rangle$. In contrast, a dressed qubit relaxes into a mixed state, symmetrizing the $|0\rangle$ and $|1\rangle$ logical states. The state readout time may exceed the qubit’s decoherence time, so having a mixed steady state helps prevent a qubit flip due to thermal relaxation, thus improving the efficiency of initial state determination. However, most quantum algorithms require pure states, making the steady state of a dressed qubit unsuitable for qubit initialization. A potential solution is to suppress the dressing amplitude $\Omega_{R}$ and apply a Zeeman field along the $z$-direction, enabling the qubit to decay into its bare ground state via thermal relaxation.
@@ Acknowledgements.
Appendix A Resonance frequency shift
If the drive amplitude (Rabi frequency) $\Omega_{R}$ is close to resonance with the harmonic oscillator, we may apply the rotating wave approximation to the qubit-harmonic oscillator to simplify the coupling term. Working in the dressed qubit frame of reference, the Hamiltonian becomes:
$\hat{H}_{\mathrm{dress}}=\frac{\Omega_{R}}{2}\hat{\sigma}_{z}+\omega_{h}\hat{a}^{\dagger}\hat{a}+g(\hat{\sigma}_{-}\hat{a}^{\dagger}+\hat{\sigma}_{+}\hat{a})\;.$ (A.1)
The coupling term in (A.1) conserves the total number of quanta of the system, and it can be exactly diagonalized in the $\{|n,1\rangle,|n+1,0\rangle\}$ subspaces [54].
Hamiltonian (A.1) can be further simplified in the dispersive regime, where the detuning between the dressed qubit and the harmonic oscillator $\Delta_{R}=\Omega_{R}-\omega_{h}$ is large compared to the coupling strength $|\Delta_{R}|\gg g$ and $\langle\hat{a}^{\dagger}\hat{a}\rangle\ll\Delta_{R}^{2}/(4g^{2})$ [19, 55]. Applying a Schrieffer-Wolff transformation [56] of the form $\hat{U}=e^{g/\Delta_{R}(\hat{a}\hat{\sigma}_{+}-\hat{a}^{\dagger}\hat{\sigma}_{-})}$ provides the Hamiltonian in the new frame,
$\hat{H}^{\prime}_{\mathrm{dress}}=\left(\frac{\Omega_{R}}{2}+\frac{g^{2}}{2\Delta_{R}}\right)\hat{\sigma}_{z}+\left(\omega_{h}+\frac{g^{2}}{\Delta_{R}}\hat{\sigma}_{z}\right)\hat{a}^{\dagger}\hat{a}.$ (A.2)
In this regime, the qubit and the harmonic oscillator do not exchange energy, but the resonance frequency of the harmonic oscillator gets shifted by ±$\pm\frac{g^{2}}{\Delta_{R}}$ depending on the state of the qubit. This shift of the resonance frequency can be used as a readout mechanism to measure the state of the qubit [28, 32, 33, 16].
Appendix B Steady state of the dressed qubit
At $T=0$, a bare qubit relaxes to its ground state due to environment-induced damping. However, this is not the case for a dressed qubit, where the qubit-heat bath interaction no longer describes energy quanta exchange with the heat bath. In this section, we derive the steady state of a dressed qubit coupled to a harmonic oscillator. Some derivations were performed using Mathematica, the corresponding notebook is available at [34]. To simplify the calculations, we adopt the rotating wave approximation taken in Sec. A. At low temperatures, the harmonic oscillator is thermalized in its ground state. The qubit-harmonic oscillator interaction [see Eq. (A.1)] couples the states $|1,n=0\rangle$ and $|0,n=1\rangle$, where the second index denotes the oscillator’s quantum number. Consequently, we can restrict our analysis to the subspace spanned by $\{|1,n=0\rangle,|0,n=0\rangle,|0,n=1\rangle\}$. In this basis, the Hamiltonian takes the form
$\hat{H}_{\mathrm{dress}}=\begin{pmatrix}\frac{\Omega_{R}}{2}&0&g\\ 0&-\frac{\Omega_{R}}{2}&0\\ g&0&-\frac{\Omega_{R}}{2}+\omega_{h}\end{pmatrix}\;.$ (B.1)
The steady state may be computed from the quantum master equation (7) by setting $\dot{\hat{\rho}}=0$. Transforming the Lindblad operators into the dressed-qubit basis and restricting to the relevant subspace, the equation reduces to a linear system for the density matrix elements, allowing for a closed-form solution.
We first examine the steady state of a dressed qubit in the absence of coupling to the harmonic oscillator $g=0$. In this case, the $|0,n=1\rangle$ state is decoupled from the states with $n=0$, meaning its associated coefficients vanish. For weak damping $\gamma_{q}\ll\Omega_{R}$, taking the $\gamma_{q}\rightarrow 0^{+}$ limit, the density matrix describing the steady state simplifies to
$\hat{\rho}_{\mathrm{st}}=\frac{1}{2}(|0\rangle\langle 0|+|1\rangle\langle 1|)\;,$ (B.2)
where we have omitted the harmonic oscillator’s quantum number for simplicity. Equation (B.2) describes a qubit in a maximally mixed state. The qubit dressing symmetrises the $|0\rangle$ and $|1\rangle$ states (and, more generally, all pure states), causing them to decay into an equidistant maximally mixed state.
Next, we analyze a dressed qubit coupled to a harmonic oscillator. In the dispersive regime $g\ll|\Delta_{R}|$. Moreover, we consider weak damping rates satisfying $\gamma_{q},\gamma_{h}\ll g$. Expanding the closed-form solution to zeroth order in the ratio of the damping rates to $g$ and subsequently to second order in $g$, the elements of the density matrix become
$\rho_{1,1}$ $=\frac{1}{2}-\frac{\gamma_{q}+4\gamma_{h}}{\tilde{\gamma}}\frac{g^{2}}{\Delta_{R}^{2}}$ (B.3a)
$\rho_{2,2}$ $=\frac{1}{2}+\frac{4\gamma_{h}-\gamma_{q}}{\tilde{\gamma}}\frac{g^{2}}{\Delta_{R}^{2}}$ (B.3b)
$\rho_{3,3}$ $=\frac{2\gamma_{q}}{\tilde{\gamma}}\frac{g^{2}}{\Delta_{R}^{2}}$ (B.3c)
$\rho_{1,3}$ $=\frac{g}{2\Delta_{R}},\;\rho_{1,2}=\rho_{2,3}=0\;.$ (B.3d)
where $\tilde{\gamma}=\frac{16\gamma_{q}\gamma_{h}}{5\gamma_{q}+4\gamma_{h}}$. As shown in Eq. (B.3), the coupling between the $|1,n=0\rangle$ and $|0,n=1\rangle$ states introduces a correction of order $O(g^{2}/\Delta_{R}^{2})$ to their population in the density matrix, decreasing $\rho_{1,1}$ and increasing $\rho_{3,3}$. These corrections reduce the qubit polarization $\langle\sigma_{z}\rangle$ ($\langle\sigma_{x}\rangle$ in the bare qubit basis used in the main text)
$\langle\sigma_{z}\rangle=-\frac{2(\gamma_{q}+4\gamma_{h})}{\tilde{\gamma}}\frac{g^{2}}{\Delta_{R}^{2}}$ (B.4)
In summary, the steady state of a dressed qubit is a maximally mixed state. However, coupling to the harmonic oscillator drives the system toward a steady state that is closer to the ground state of the dressed qubit.
Appendix C Quantum bus between two dressed qubits
In this section, we consider two driven qubits coupled to a harmonic oscillator. The harmonic oscillator works as a quantum bus between qubits, entangling both qubits [19]. The Hamiltonian for the system can readily be obtained by generalizing Eq. (A.1),
$\hat{H}_{\mathrm{dress}}=\sum_{j}\frac{\Omega_{R,j}}{2}\hat{\sigma}_{z}^{j}+\omega_{h}\hat{a}^{\dagger}\hat{a}+\sum_{j}g_{j}(\hat{\sigma}_{-}^{j}\hat{a}^{\dagger}+\hat{\sigma}_{+}^{j}\hat{a})\;,$ (C.1)
where $j=1,2$ is the qubit index. Following the same steps as in Appendix A, applying a Schrieffer-Wolff transformation of the form $\hat{U}=e^{\sum_{j}g_{j}/\Delta_{R,j}(\hat{a}\hat{\sigma}_{+}^{j}-\hat{a}^{\dagger}\hat{\sigma}_{-}^{j})}$, in the dispersive regime the Hamiltonian acquires an effective qubit-qubit interaction term:
$\hat{H}^{\prime}_{\mathrm{dress}}=\sum_{j}\left(\frac{\Omega_{R,j}}{2}+\frac{g_{j}^{2}}{2\Delta_{R,j}}\right)$ $\hat{\sigma}_{z}^{j}+\left(\omega_{h}+\sum_{j}\frac{g_{j}^{2}}{\Delta_{R,j}}\hat{\sigma}_{z}^{j}\right)\hat{a}^{\dagger}\hat{a}$ (C.2)
$+\sum_{j}\frac{g_{j}g_{\bar{j}}}{2\Delta_{R,j}}(\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{\bar{j}}+\hat{\sigma}_{-}^{j}\hat{\sigma}_{+}^{\bar{j}})\;,$
where $\bar{j}$ is the qubit opposite to $j$.
If the transition frequencies and the coupling constants of both qubits are the same, the interaction term (second line of Eq. (C.2)) commutes with the noninteracting Hamiltonian part $\hat{H}_{0}$, so the time evolution operator is given by [37, 57, 58]
[…]
If the qubit-qubit interaction time is fixed to $t=\frac{\pi|\Delta_{R}|}{4g^{2}}$ and the harmonic oscillator remains in the ground state, $\hat{U}(t)$ corresponds to a $\sqrt{\mp i\mathrm{SWAP}}$-gate depending on the sign of the detuning $\Delta_{R}$ [see Eq. (11)], up to 1-qubit gates. The interaction time can be controlled by tuning the drive amplitudes $\Omega_{R,j}(t)$ of the qubits, such that an appropriate pulse length tunes the dressed qubits in resonance with the harmonic oscillator. The $\sqrt{i\mathrm{SWAP}}$-gate is equivalent to a CNOT gate up to 1-qubit gates, so the interaction in Hamiltonian (C.2) is sufficient for universal quantum computation [5].
Figure S.5: Generation of a Bell state in the computational basis.
Appendix D Creating Bell states with dressed qubits
[…]